### v1

# AECBench: A Hierarchical Benchmark for Knowledge Evaluation of Large Language Models in the AEC Field

Abstract
Large language models (LLMs), as a novel information technology, are seeing increasing adoption in the Architecture, Engineering, and Construction (AEC) field. They have shown their potential to streamline processes throughout the building lifecycle. However, the robustness and reliability of LLMs in such a specialized and safety-critical domain remain to be evaluated. To address this challenge, this paper establishes AECBench, a comprehensive benchmark designed to quantify the strengths and limitations of current LLMs in the AEC domain. The benchmark features a five-level, cognition-oriented evaluation framework (i.e., Knowledge Memorization, Knowledge Understanding, Knowledge Reasoning, Knowledge Calculation, and Knowledge Application). Based on the framework, 23 representative evaluation tasks were defined. These tasks were derived from authentic AEC practice, with scope ranging from codes retrieval to specialized documents generation. Subsequently, a 4,800-question dataset encompassing diverse formats, including open-ended questions, was crafted primarily by engineers and validated through a two-round expert review. Furthermore, an “LLM-as-a-Judge” approach was introduced to provide a scalable and consistent methodology for evaluating complex, long-form responses leveraging expert-derived rubrics. Through the evaluation of nine LLMs, a clear performance decline across five cognitive levels was revealed. Despite demonstrating proficiency in foundational tasks at the Knowledge Memorization and Understanding levels, the models showed significant performance deficits, particularly in interpreting knowledge from tables in building codes, executing complex reasoning and calculation, and generating domain-specific documents. Consequently, this study lays the groundwork for future research and development aimed at the robust and reliable integration of LLMs into safety-critical engineering practices.
11footnotetext: Corresponding author: Xianzhong Zhao (x.zhao@tongji.edu.cn)22footnotetext: Corresponding author: Ruifeng Luo (ruifengluo.tj@gmail.com)
Keywords Large Language Models; Domain-specific Benchmark; Architecture, Engineering, and Construction (AEC); Knowledge Evaluation; Automated Evaluation
1 Introduction
With the ongoing intelligent transformation of the Architecture, Engineering, and Construction (AEC) industry, information technologies have played a crucial role in advancing engineering practices [1]. Among these technologies, large language models (LLMs) hold the potential to augment the knowledge-intensive tasks inherent in a building’s lifecycle, leveraging their vast knowledge repositories and cross-domain application capabilities [2, 3, 4]. To effectively harness this potential and guide the future development of LLMs for the industry, a standardized and domain-specific evaluation benchmark is essential for assessing LLMs’ proficiency in these contexts [5, 6, 7, 8].
The highly specialized and safety-critical nature of the AEC domain presents a unique challenge for LLMs. As an interdisciplinary convergence of civil engineering and allied fields, the AEC domain demands extensive domain knowledge and operational capabilities spanning the entire lifecycle of a project. Figure 1 illustrates the multifaceted knowledge required in this integrative paradigm, encompassing 11 key domains, as well as numerous specialized topics radiating from these central domains. These domains are highly interrelated and collectively contribute to successful project completion. For instance, structural design is not solely governed by the principles of mechanics; it must also integrate considerations of architectural aesthetics, material properties, construction techniques, fire safety, energy efficiency, and environmental regulations, among numerous other codes and standards. Given that the AEC domain is safety-critical, this interconnectivity places stringent demands on the accuracy and reliability of LLMs. These models must be capable of assisting engineers by retrieving relevant information, facilitating cross-disciplinary decisions, and providing verifiable, evidence-based recommendations.
Recent studies have demonstrated the rapid integration of LLMs across the lifecycle of the AEC industry. In the design phase, LLMs are used to automate the generation of building energy models (BEMs) from descriptive text [9] and to enable intuitive, conversational querying of building information models (BIMs) [10]. Furthermore, LLMs are applied to automate regulatory compliance by converting textual building codes into machine-executable rules [11, 12, 13]. In structural engineering, LLMs function as controllers for the generative design of structural elements such as shear walls [14]. LLMs also facilitate structural health monitoring by analyzing sensor data [15]. During the construction phase, applications range from automating management tasks like project scheduling [16] to significantly enhancing site safety. Safety improvements include increasing the reliability of safety information retrieval by comparing retrieval-augmented generation (RAG) with fine-tuning techniques [17, 18], automatically analyzing accident reports to identify root causes [19], and developing interactive, game-based training platforms to improve worker safety knowledge [20]. These studies exemplify the potential of LLMs within the domain-specific tasks. However, they often focus on specific, narrow scenarios, leaving their broader capabilities untested. Consequently, their findings are difficult to generalize, which impedes cross-study comparability.
Advancing the robustness and reliability of LLMs in the AEC domain requires the establishment of domain-specific benchmarks. These benchmarks provide quantitative metrics to evaluate a model’s grasp of domain knowledge and its performance on relevant tasks. They play a crucial role throughout the LLM lifecycle. For practitioners, these benchmarks serve as a vital tool for model selection, offering a reliable, evidence-based means to identify the LLM best suited for their specific application needs [5, 7, 8]. For LLM developers and researchers, they provide a standardized and quantitative methodology to assess model performance, identify areas for improvement, and drive the innovation of more capable and specialized models [5, 21, 8]. Finally, for regulators, they offer the necessary empirical data to inform the creation of robust safety standards, ethical guidelines, and governance frameworks, ensuring the responsible deployment of this novel information technology in high-stakes fields [6].
Domain-specific benchmarks are being actively developed in various fields, including law, healthcare, and finance. LawBench [22] categorizes LLMs’ capabilities in law into three cognitive levels (i.e., legal knowledge memorization, understanding, and applying) and provides detailed tasks and questions for each. PromptCBLUE [23] adapts various tasks in CBLUE [24] (for example, medical named entity recognition and event extraction) into a unified language generation format to set a standard for LLMs in Chinese medical contexts. FinEVAL [25] uses finance-related multiple-choice questions to measure the effectiveness of LLMs in finance. Although benchmark frameworks developed for these fields provide valuable methodologies, they cannot be directly applied to the AEC field due to significant differences in domain characteristics. The AEC domain encompasses multidisciplinary knowledge, and the tasks require not only legal-analogous capabilities in code comprehension and logical reasoning (e.g., building code interpretation), but also finance-inspired calculation proficiency (e.g., structural load calculation), as well as unique domain-specific capabilities (e.g., structural stress analysis).
Several benchmarks have emerged to evaluate LLMs’ performance in the AEC domain. Qin et al. curated a benchmark comprising 100 fill-in-the-blank and true/false questions derived from textbook content [26], specifically from the topic "Basic Principles of Concrete Structures" within the domain of "Principles of Structural Design". Wu et al. created a larger benchmark of 875 multiple-choice questions covering a wide range of topics in the architectural domain, encompassing design, engineering, and planning knowledge [27]. These benchmarks offer valuable initial contributions to the evaluation. However, they share the limitation of a relatively narrow evaluation scope. Compared to the scope of knowledge shown in Figure 1, Qin’s benchmark covers just a part of one topic: Design of Reinforced Concrete and Steel Structures. Similarly, Wu’s benchmark is confined to the single key domain of Principles of Architectural Design. Liang et al. developed AEC-Bench, a multiple-choice evaluation consisting of over 1,000 questions sourced from official licensure examinations. This benchmark assesses abilities across twelve disciplines and five cognitive levels: knowledge memorization, understanding, reasoning, simple computation, and complex calculation [28]. Their findings reveal a progressive performance decline from lower-order cognitive tasks (memorization and understanding) to higher-order ones (reasoning and calculation). However, a key limitation of aice.AEC-Bench is its exclusive reliance on multiple-choice questions from examinations. Such questions diverge considerably from real-world AEC scenarios, where professionals are frequently required to address open-ended and ill-structured problems.
Therefore, existing evaluation benchmarks are fundamentally limited on two fronts. First, they provide an inadequate assessment of the breadth and depth of domain knowledge. Second, their misalignment with real-world scenarios means they fail to offer meaningful or trustworthy insights into a model’s practical skills. In view of these limitations, this paper presents AECBench, a comprehensive Chinese benchmark specifically designed to evaluate LLMs in real-world AEC scenarios across diverse cognitive levels. The main innovations are as follows:
  • Hierarchical Cognitive Framework: A novel hierarchical cognitive framework tailored for the AEC domain is introduced. This framework is designed to mirror the graduated level of cognitive demands found in AEC tasks, from basic knowledge memorization to complex application. The framework is structured into five distinct levels. Notably, it includes dedicated levels for Knowledge Reasoning and Knowledge Calculation, which are two critical skills in the AEC domain. For a more fine-grained assessment of practical abilities, the Knowledge Application level is further partitioned into three sub-levels: Analysis, Evaluation, and Creation. A more comprehensive and realistic evaluation of an LLM’s capabilities is made possible by this hierarchical design, through which its strengths and weaknesses across different cognitive levels can be identified.
  • High-Quality Benchmark Dataset: A high-quality benchmark dataset consisting of 4,800 meticulously crafted or curated questions across 23 evaluation tasks was constructed, characterized by three core strengths. First, the evaluation tasks were curated by domain engineers to reflect real-world scenarios in the AEC field, with questions independently crafted or curated from building codes, proprietary internal documents, and licensure examinations. Second, the dataset offers a diverse range of question formats, including multiple-choice, classification, extraction, and open-ended generation, designed to authentically reflect the complexity and dynamic, open-ended nature of real-world AEC scenarios and professional workflows. Third, each question was subjected to multiple rounds of data review by top professionals from leading AEC enterprises, ensuring high clarity and accuracy throughout the dataset.
  • Automated Evaluation Pipeline and Open-source Release: An automated evaluation pipeline for open-ended, long-form responses is introduced. It employs an "LLM-as-a-judge" approach to score documents against the rubrics established by domain experts. This method significantly enhances evaluation scalability and speed while ensuring that the assessment remains grounded in expert knowledge. The entire AECBench, including the code and dataset, is to be released as an open-source resource for the community.
2 The design methodology of AECBench
2.1 Hierarchical evaluation framework
This section aims to comprehensively and effectively assess the capabilities of large language models by reflecting the unique knowledge, application scenarios, and cognitive challenges inherent to the AEC domain. Challenging the conventional approach of categorizing tasks by their disciplines or difficulties [7, 26, 27], this paper argues for a more nuanced evaluation framework for the complex AEC field. The rationale is twofold: (i) performance is dictated more by a task’s cognitive demands than its subject matter, and (ii) AEC tasks themselves are rarely cognitively monolithic. For example, a task in "Design of Reinforced Concrete and Steel Structures" necessitates memorizing code provisions, understanding load-bearing principles, reasoning about structural integrity, and creating a compliant design document. To adequately capture this complexity, this paper establishes a hierarchical framework grounded in these distinct cognitive abilities, referencing Bloom’s Taxonomy. Bloom’s Taxonomy, first introduced in 1956 and revised in 2001, is a well-established cognitive taxonomy [29, 30]. It categorizes learning objectives into six progressive levels: Remember, Understand, Apply, Analyze, Evaluate, and Create. Due to its hierarchical structure, which maps a progression from foundational knowledge to advanced problem-solving, it provides a robust theoretical basis for the proposed evaluation framework.
The hierarchical evaluation framework consists of five distinct levels: Knowledge Memorization, Knowledge Understanding, Knowledge Reasoning, Knowledge Calculation, and Knowledge Application (see Figure 1). The proposed framework adapts and extends Bloom’s Taxonomy to cater to the specific demands of the AEC domain. Specifically, it retains the foundational levels of Remember and Understand (i.e., Knowledge Memorization and Knowledge Understanding), while grouping the higher-order skills, Analyze, Evaluate, and Create, as distinct sub-levels within a comprehensive Knowledge Application level. Furthermore, to address the unique requirements of the field, two dedicated levels are introduced: Knowledge Reasoning and Knowledge Calculation. This extension is motivated by the prevalence of tasks requiring logical deduction or numerical computation across key AEC domains (as shown in Figure 1), a necessity that is particularly prominent in areas such as mechanics, principles of structural design, and engineering economics.
Figure 1: The hierarchical evaluation framework
The Knowledge Memorization level assesses the model’s ability to recall foundational AEC facts. The Knowledge Understanding level evaluates LLMs’ capacity to not only retain but also meaningfully interpret and contextualize foundational AEC knowledge, bridging memorized facts with their practical significance and logical connections. The Knowledge Reasoning level assesses an LLM’s ability to utilize its internalized domain knowledge, synthesizing multiple facts, codes, and contextual factors to reach sound, defensible conclusions. This level assesses the model’s ability to decompose intricate problems as well as to identify and reconcile relevant constraints. Furthermore, it evaluates whether the model can construct a clear, step-by-step logical chain that leads to viable decisions and judgments. The Knowledge Calculation level assesses an LLM’s ability to apply AEC knowledge to produce verifiable quantitative results. It evaluates whether the LLM can select appropriate formulas, use correct parameters, maintain dimensional consistency, and carry out multi-step calculations. The Knowledge Application level is uniquely tailored to the applied nature of the AEC discipline, and structured into three distinct sub-levels: Analysis, Evaluation, and Creation. Progressing through three ascending levels, it encompasses: Analysis, in which the model inductively extracts, classifies, or verifies domain-specific data or information; Evaluation, where it assesses whether a given document or solution meets the AEC sector’s requirements; and Creation, where it synthesizes and produces domain-specific documents such as design proposals, technical specifications, or construction schedules that are both novel and technically sound.
Notably, the Knowledge Reasoning and Knowledge Calculation levels assess abilities that share conceptual similarities with those evaluated in other major benchmarks. For instance, these abilities are similar to the mathematical reasoning skills tested in prominent benchmarks like GSM8K [31] and Math-500 [32]. These conceptual similarities are also evident when compared to domain-specific benchmarks such as FinanceMath [33], MedCalc-Bench [34], and LawBench [22] for the financial, medical, and legal domains, respectively. This conceptual alignment with established benchmarks underscores the framework’s focus on critically recognized LLM competencies.
2.2 Task design
A total of 23 tasks were involved in the evaluation framework for the various cognitive levels (see Figure 2). These tasks were all derived from real-world AEC scenarios and were curated by domain engineers.
Figure 2: Designed evaluation tasks of AECBench
2.2.1 Knowledge Memorization tasks
This paper establishes three tasks designed to assess the memorization of key AEC knowledge, specifically targeting three typical types: (i) core law articles and code provisions, (ii) fundamental terminology, and (iii) domain-specific abbreviations.
  • Code Memorization & Retrieval (1-1): Given an exact excerpt from a provision of building codes, identify the originating code by selecting the correct option from four candidates.
  • Terminology Memorization (1-2): Given a definition of a domain-specific term, identify the correct terminology corresponding to the definition by selecting the correct option from four candidates.
  • Abbreviation Memorization (1-3): Given a domain-specific abbreviation, identify its correct expansion by selecting the correct option from four candidates.
2.2.2 Knowledge Understanding tasks
To assess the comprehension of knowledge from two critical sources, this paper establishes two types of tasks: one focused on standards and codes, and the other on established design literacy.
  • Code Provision Interpretation (2-1): Given a partially paraphrased description related to a code provision, the model must demonstrate comprehension of the provision’s qualitative or quantitative requirements by selecting the correct option from four candidates.
  • Design Literacy Q&A (2-2): Given an incomplete exposition of domain design literacy, the model must infer the missing core concept and select the correct option from four candidates.
  • Code Provision Interpretation(Tabular data)(2-3): Given a partial description related to a table in the code provision, the model must demonstrate comprehension of the provision’s qualitative or quantitative requirements by selecting the correct option from four candidates.
It is noted that the code provision interpretation task is further divided into two sub-tasks, and a task (Task 2-3) that requires the model to comprehend knowledge contained within code tables is designed. Figure 3(a) illustrates an example of the question generation process from tabular data. In the figure, a multiple-choice question was crafted based on Table 6.4.5-1 of "Code for seismic design of buildings" [35], requiring the model to select the maximum axial compression ratio for a seismic wall under specific conditions.
Figure 3: Illustrative examples of the two proposed tasks
(a) the code provision interpretation (tabular data) task; (b) the design proposal generation task
2.2.3 Knowledge Reasoning tasks
This level includes two representative tasks designed to evaluate LLMs’ logical reasoning abilities: a deductive task that requires making design decisions based on established knowledge, and an inductive task that involves inferring a building’s type from its description.
  • Design Decision Formulation (3-1): Given a descriptive scenario from a project (e.g., project specifications, site conditions), employ deductive reasoning to determine which of four candidates, representing possible measures, action decisions, or patterns of change, is both logically coherent and compliant with the applicable codes and standards.
  • Building Type Inference (3-2): Given a descriptive passage about a specific building, apply inductive reasoning to infer its building type and select the correct option from four candidates.
  • Design Decision Formulation (Tabular Data) (3-3): This task is identical to Task 3-1, except that it requires referencing knowledge contained in the code tables.
Similar to the task design process in the Knowledge Understanding level, the design decision formulation task is further divided into two sub-tasks based on whether tabular code knowledge is required.
2.2.4 Knowledge Calculation tasks
Four typical calculation tasks are involved in this level, each designed to assess quantitative abilities at a key stage of the building lifecycle: architectural and structural design, MEP design, construction, and engineering economics.
  • Calculation in Architectural Design (4-1): Given a quantitatively formulated problem in the domain of architectural and structural design (e.g., calculating the moment of inertia of a cross-section), perform the requisite numerical analyses to determine the single correct solution from four candidates.
  • Calculation in MEP Design (4-2): This task is identical to Task 4-1, but it focuses on assessing calculation abilities in MEP design (e.g., calculating the airflow rate of a fan).
  • Calculation in Construction (4-3): This task is identical to Task 4-1, but it focuses on assessing calculation abilities in construction (e.g., calculating the project duration and worker shift scheduling).
  • Calculation in Engineering Economics (4-4): This task is identical to Task 4-1, but it focuses on assessing calculation abilities in engineering economics (e.g., calculating the interest and net present value).
2.2.5 Knowledge Application tasks
At the Knowledge Application level, multiple representative tasks were designed for each sub-level. Five representative application tasks are considered at the Analysis sub-level: document classification, document proofreading, compliance checking, brand compliance verification, and extraction of project information. At the Evaluation sub-level, this paper primarily establishes tasks on assessing the LLM’s capabilities in acting as a rubric-based judge to score documents in architectural and structural design. At the Creation sub-level, this paper designs tasks that require LLMs to generate three typical types of documents in domain-specific workflows: conceptual design proposals, specialized reports, and tender review reports.
Analysis
  • Document Classification (5-1-1): Given a sentence extracted from a domain-specific document, determine and assign the sentence to its appropriate professional discipline category.
  • Document Proofreading (5-1-2): Given a sentence drawn from a domain-specific document, identify and rectify all linguistic defects, including spelling mistakes, redundancies, omissions, word-order errors, and semantic inaccuracies, and produce a coherent, grammatically correct, and semantically faithful revision.
  • Compliance Checking (5-1-3): Given a design case with multiple parameters, the model is required to analyze whether the design complies with code provisions through calculations and logical reasoning. If non-compliant, the model must identify the specific cause of the violation. Finally, select the correct violation reason from four candidates.
  • Brand Compliance Verification (5-1-4): Given a tabular equipment specification sheet, detect and report all potential brand-related deviations present in the corresponding bid submissions.
  • Information Extraction (5-1-5): Given a long, domain-specific document (e.g., a tender announcement), identify and extract every entity mention that matches a predefined inventory of entity types.
Evaluation
  • Evaluation of Documents in Architectural Design (5-2-1): Given a design scenario and an architectural proposal, the model is required to act as a judge and score the proposal based on predefined criteria, including content accuracy, professional expression, readability, and formatting.
  • Evaluation of Documents in Structural Design (5-2-2): This task is identical to Task 5-2-1, but it focuses on evaluating documents in structural design.
Creation
  • Conceptual Design Proposal Generation (5-3-1): Given a project brief and its associated design specifications, automatically generate a preliminary design proposal.
  • Specialized Report Generation (5-3-2): Given the essential project specifications, automatically generate a structural design basis report that adheres to established codes.
  • Tender Review Report Generation (5-3-3): Given a device specification file and an equipment specification table, the model must first identify all parameter mismatches and then generate a coherent tender review report that summarizes these discrepancies.
2.3 Question design
2.3.1 Diverse question formats in AECBench
AECBench employs a diverse set of question formats including multiple-choice, classification, extraction, and generation. Figure 2 shows the number of questions corresponding to each individual task. Multiple-choice questions (MCQs) were used across the Knowledge Memorization, Understanding, Reasoning, and Calculation levels, as well as for the compliance checking task within the Knowledge Application level. They offer clear evaluation metrics (i.e., accuracy) and serve as a simple but effective proxy for assessing advanced capabilities of LLMs. For all tasks within the Knowledge Application level, excluding the compliance checking task, the question formats were intentionally diversified and specifically tailored to the requirements of each task. The classification format was applied to the document classification task; the extraction format was applied to the information extraction task; and a generative format was applied to both the document proofreading and brand compliance verification tasks. It must be mentioned that open-ended questions were utilized for the tasks at the Creation sub-level. Details associated with the open-ended questions are presented in Section 2.3.2.
By moving beyond a sole reliance on the multiple-choice format, the benchmark mitigates inherent limitations such as option-setting bias and the possibility of correct guessing [36]. This approach also enables a more robust evaluation of a model’s abilities in authentic AEC practice, which are difficult to assess with multiple-choice questions [37, 38].
2.3.2 Open-ended questions and automated rubric-based evaluation
The proposed AECBench highlights the use of open-ended questions to rigorously assess a model’s ability to produce complex and lengthy engineering documents. Instead of simply generating a "yes/no" or predetermined response, the model is required to generate professional documents (for instance, conceptual design proposals) with approximately 2,000 words following an instruction and a prompt. Open-ended questions provide deeper insight into the model’s generative capabilities, nuanced contextual understanding, and ability to tackle complex tasks.
A rubric-based assessment pipeline is established to facilitate the rapid and scalable evaluation of open-ended questions, where LLMs act as automated evaluation agents. This pipeline follows a two-stage workflow. Firstly, the domain experts predefine a detailed set of rubric criteria that serve as the evaluation reference. Subsequently, using the LLM-as-a-Judge approach [39], a generated document is assessed by scoring it against these expert-defined rubrics.
Figure 3(b) presents an example of the question and its rubric in the specialized report generation task. The example in the figure requires determination of the seismic design parameters for a specific building. The rubric breaks down the problem into five sequential steps, each of which is awarded partial credit, to assess the model’s ability to apply codes. This design shifts expert effort from exhaustive grading to the upfront definition of rubrics, thereby leveraging crucial expert knowledge, lowering human workload, and greatly enhancing both scalability and objectivity.
To validate the reliability of the automated evaluation pipeline, this paper establishes parallel validation tasks in the Evaluation sub-level (Tasks 5-2-1 and 5-2-2). This involves tasks where both human experts and LLMs act as judges to score the same documents. A schematic overview of Evaluation and Creation tasks assessment methods is presented in Figure 4. In the Evaluation tasks, the model exhibiting the highest correlation with the scores of human experts is selected and subsequently employed as the judge for the generation tasks.
Figure 4: Automated evaluation pipeline for open-ended questions
3 Dataset construction
A custom dataset comprising 4,800 questions was constructed based on the evaluation tasks considered in the hierarchical cognitive framework. The dataset was established through three steps: data collection, data cleaning, and data review (see Figure 5). The data review process highlights a two-round review mechanism (i.e., each individual item is reviewed by a mid-level engineer and then confirmed by an expert), ensuring that a high-quality dataset for LLM evaluation is obtained.
Figure 5: The workflow of constructing the dataset
3.1 Data collection
The data sources for AECBench evaluation dataset are twofold, with a primary emphasis on questions meticulously crafted by a cross-disciplinary team of domain engineers. Guided by the predefined evaluation framework, the engineers developed targeted assessment items by systematically integrating building codes, proprietary internal materials, and other domain-specific documents, ensuring precise alignment with the cognitive hierarchies and the designed tasks. To ensure foundational coverage, items from official engineering licensure examinations were also included. These questions, sourced from examinations for registered architects, design engineers, and constructors, are valuable for practitioners and address key knowledge across the entire project lifecycle.
3.2 Data cleaning
The data cleaning process begins with converting all raw data into a standardized format. This is followed by a multi-step procedure including: question deduplication, which uses n-gram overlap and vector similarity algorithms to remove redundant questions; outlier detection, which focuses on correcting factual and spelling anomalies; and content completion, which verifies the completeness of answers and options. This rigorous process ensures the dataset is reliable and ready for data review.
3.3 Data review
The data review highlights a two-round process: preliminary content review and expert review. The preliminary content review is performed by mid-level engineers, who evaluate the reasonableness of each question’s content and ensure the overall dataset covers key industry topics. This serves as an effective first line of defense against straightforward inaccuracies. The subsequent stage is the expert review, where senior specialists (engineers with over 10 years of experience) apply their deep expertise to assess the data’s representativeness and identify subtle content errors. This expert-led process serves as the core quality control protocol and was found to be critical for identifying errors, clarifying ambiguities, and enhancing overall data reliability.
Figure 5 visually details the necessity of this two-round data review mechanism through a specific example. In the initial data collection phase, question-and-answer pairs concerning the seismic design for buildings were generated. This review was primarily tasked with checking and correcting foundational, factual errors. As shown with Question Q1, an incorrect statement regarding the seismic requirements for Category C buildings was identified as being in direct contradiction with the building codes. Consequently, Question Q1 was flagged as "Fail". This initial stage served as an effective first line of defense against straightforward inaccuracies. However, a reliance on the preliminary content review alone was determined to be insufficient, as more nuanced issues require deep domain knowledge to be detected. This highlights the decisive importance of the second review stage: expert review. The criticality of this stage is demonstrated by the case of Question Q2. This question was initially assessed as "Pass" during the preliminary content review, as no overt errors were present. However, during the expert review, a critical ambiguity was identified regarding whether a "university" should be classified as a "school" under different interpretations of the code, which affects the required seismic grade. This ambiguity, which could have severely impacted question clarity, was not only identified but also resolved by the domain expert. The question was revised to specify a "primary school" to eliminate any interpretive ambiguity. As a general statement, the two-round iterative mechanism, particularly the inclusion of the second expert-led stage, significantly enhanced the reliability of the final dataset, thereby providing a robust foundation for the ensuing analysis.
4 Evaluation of large language models (LLMs)
4.1 Models
Nine widely used LLMs were selected, encompassing both open-source and proprietary architectures. with QwQ-32B, DeepSeek-R1, and GPT o3-mini featuring specialized optimizations for enhanced reasoning capabilities. The selected models and their information are shown in Table 1.
Table 1: Overview of evaluated models
Models Creator Parameters Reasoning
Moonshot-v1-128k [40] Moonshot AI Undisclosed w/o
GLM-4-Plus [41] Zhipu Undisclosed w/o
Qwen-Turbo [42] Alibaba Undisclosed w/o
QwQ-32B [43] Alibaba 32B w/
DeepSeek-V3 [44] DeepSeek 671B w/o
DeepSeek-R1 [45] DeepSeek 671B w/
GPT-4o [46] OpenAI Undisclosed w/o
GPT o3-mini [47] OpenAI Undisclosed w/
Hunyuan-TurboS [48] Tencent Undisclosed w/o
  • “w/” (with) or “w/o” (without) reasoning indicates whether the model is able to output its chain-of-thought reasoning or to give only the final answer without exposing intermediate reasoning steps .
4.2 Experiment settings
The evaluation experiments were conducted on the OpenCompass platform [49], which enables automated answer extraction from model outputs and standardized computation of benchmarking metrics through its modular architecture. The experiments were conducted under a one-shot setting, which involves evaluating the model with a single example provided for each task, as shown in Figure 6. In this case, the model is expected to understand the task and generalize from this single example with minimal guidance. All experiments were primarily carried out on a server featuring an NVIDIA 4090 GPU (24GB) and Linux Ubuntu 22.04.3 LTS.
Figure 6: One-shot example for multiple-choice questions
4.3 Main results
4.3.1 General observations
A systematic evaluation of nine LLMs was conducted through the evaluation platform. As detailed in Table 2, the task-specific performance metrics for each LLM are visualized in Figures 7 and 8. These figures also include horizontal lines representing the average performance for each level and sub-level. The test results clearly suggest a decline in LLMs’ capabilities across the five cognitive levels. A similar downward trend is also evident across the three sub-levels of the Knowledge Application level. Both major inter-level gaps and fine-grained intra-level differences were successfully identified, thereby validating the soundness and necessity of the hierarchical evaluation framework.
Table 2: Evaluation results
Task ID Moonshot-v1 -128k GLM-4 -Plus Qwen -Turbo QwQ -32B DeepSeek -V3 DeepSeek -R1 GPT -4o GPT o3-mini Hunyuan -TurboS
1-1 67.27 73.64 70.45 69.54 83.64 87.27 67.27 63.64 82.73
1-2 96 97.5 98.5 97 98 98 97.5 95 98.5
1-3 71.67 80.00 67.78 77.22 83.89 87.22 78.33 86.11 75.00
2-1 55.24 67.67 73.41 67.73 78.34 84.02 60.05 54.06 77.03
2-2 71.52 85.45 87.88 88.48 88.48 89.70 80.00 70.91 90.30
2-3 43.92 34.92 43.92 40.74 52.38 63.49 44.44 39.68 43.92
3-1 53.69 66.96 74.63 71.39 75.52 82.60 60.77 60.47 74.63
3-2 85.00 85.00 76.00 83.00 90.00 91.00 85.50 82.50 88.00
3-3 34.65 35.83 43.31 42.13 32.68 53.94 39.76 42.91 44.49
4-1 35.45 47.62 41.27 43.92 57.67 78.84 39.68 51.32 33.86
4-2 34.62 73.08 69.23 88.46 88.46 96.15 65.38 92.31 41.53
4-3 44.26 81.97 65.57 83.61 86.89 85.25 59.02 75.41 48.08
4-4 35.59 54.24 50.85 25.42 55.93 66.10 50.85 44.07 67.21
5-1-1 85.28 85.56 77.22 76.96 86.94 87.78 89.72 87.22 85.00
5-1-2 92.23 93.24 93.75 86.12 90.19 62.73 93.14 89.51 81.20
5-1-3 32.95 7.66 38.70 45.21 34.87 52.49 42.91 48.66 45.21
5-1-4 26.90 30.38 28.45 18.80 30.66 29.73 31.09 29.95 37.00
5-1-5 73.21 64.15 67.35 44.10 73.10 70.48 68.43 69.12 73.10
5-2-1 48.4 47.1 47.4 38.3 45.2 53.2 46.5 50.5 54.9
5-2-2 30.9 33.9 41.7 51.1 55.7 56.4 34.2 46.6 49.6
5-3-1 19.83 25.67 24.11 32.67 36.33 50.28* 19.33 29.11 43.94
5-3-2 27.78 36.67 33.33 30.00 39.56 44.44* 24.22 15.56 40.00
5-3-3 39.44 49.17 33.61 52.78 49.17 51.11* 44.44 53.33 42.78
  • The table shows the evaluation results of nine LLMs across twenty-three tasks. All multiple-choice questions are evaluated using accuracy as the metric. For the classification task (Task 5-1-1), the evaluation metric is F1. For Tasks 5-1-4 and 5-1-5, Soft-F1 is used as the evaluation metric, and F0.5 is used as the evaluation metric for Task 5-1-2. For evaluation tasks (Tasks 5-2-1 and 5-2-2), Kendall rank correlation coefficient ($\tau$) is used as the evaluation metric. Kendall’s $\tau$ is a statistic used to measure the ordinal association between two measured quantities. It ranges from -1 to +1, where +1 indicates perfect agreement between the two rankings, -1 indicates perfect disagreement, and a value near 0 indicates no correlation. In Tasks 5-3-1, 5-3-2, and 5-3-3, given that the DeepSeek-R1 model serves as the expert model for benchmark scoring, all corresponding performance metrics are annotated with an asterisk (*) to alert readers to potential evaluation bias risks. To provide a unified and comparable view of performance across all tasks, the results for all metrics have been normalized to a 0-100 scale. As all observed Kendall’s $\tau$ values were positive ([0, 1]), they were scaled by 100 (Score = $\tau$ × 100) for clarity and consistent presentation. The numbers in bold in the table represent the best-performing model for each task.
A noticeable performance drop was identified between the Knowledge Memorization and Understanding levels, which is attributed to the cognitive leap required to move beyond the retrieval of explicitly encoded facts to the construction of deeper logical relationships. Another steep decline in performance was evident among the sub-levels of knowledge application, particularly in the creation sub-level. This decline is compounded by a methodological factor, as a transition is made from a uniform multiple-choice format to diverse task types with varied evaluation metrics.
Figure 7: Performance of models in Knowledge Memorization/Understanding/Reasoning/Calculation tasks
Figure 9 highlights a critical disparity between the model’s knowledge memorization and its practical application. The model demonstrated this gap in complex creative tasks; for instance, it consistently failed to determine the correct seismic grade when generating a full design proposal. However, the model could accurately provide the very same seismic grade when asked for it in a direct, isolated query. This proves the model’s failure is not a knowledge deficit but an application deficit. While the model may possess the necessary factual knowledge, it cannot reliably access and integrate it into a multi-step reasoning and long-form generation process.
Figure 8: Performance of models in Knowledge Application tasks
The DeepSeek-R1 demonstrates a significant performance advantage over DeepSeek-V3, excelling in 19 of 23 designed tasks. This superiority is most pronounced in tasks at the Knowledge Reasoning and Knowledge Calculation level and the Evaluation and Creation sub-level. A direct comparison between the two models is justified by their virtually identical model architectures and parameter scales. This finding suggests that, for models of equivalent scale and architecture, reinforced training that specifically targets reasoning capabilities yields substantial enhancements in handling complex tasks within the AEC sector. For context, other control pairs from the same organizations, such as Qwen-32B vs. Qwen-Turbo and GPT-4o vs. GPT-3-mini, were considered. However, a lack of publicly available data on their internal architectures and parameters precludes a meaningful comparison, and their performance benchmarks do not present a discernible trend.
Among all the models, DeepSeek-R1 exhibits superior performance, standing out as the best-performing model in 14 out of 23 tasks and is validated as the automated judge. However, despite its strong and broad capabilities, its performance underscores a broader finding: even the most advanced LLMs still have substantial room for improvement in handling complex AEC-related tasks.
Figure 9: Failure case analysis: the gap between Knowledge Memorization and Application
4.3.2 Specific observations
At the Knowledge Memorization level, all models performed exceptionally well, with scores universally exceeding 95% on terminology memorization (Task 1-2). This strongly suggests that a considerable amount of foundational AEC knowledge is already embedded within the parameters of current LLMs.
At the Knowledge Understanding and Reasoning levels, a notable weakness was observed in tasks requiring the interpretation of tabular data (Tasks 2-3 and 3-3). The performance of the LLMs was notably deficient, even falling below that of the supposedly more advanced application tasks (for example, information extraction tasks). Conversely, on the building type inference task (Task 3-3), performance unexpectedly approached the performance of some Knowledge Memorization tasks (for example, code memorization & retrieval task). An analysis of Task 3-3 indicates that its correct option often exhibits high semantic similarity with the question stem, primarily because of shared keywords. Consequently, successfully completing this task may not strictly require inductive reasoning.
At the Knowledge Calculation level, the lower scores in architectural design calculation (Task 4-1) and engineering economics calculation (Task 4-4) are attributable to the multi-step analysis and complex computations these problems require, posing a greater challenge to the models.
At the Knowledge Application level, performance on the sub-levels was highly variable. In the Analysis tasks, models worked well on document classification and proofreading tasks (Tasks 5-1-1 & 5-1-2) but struggled notably with compliance checking (Task 5-1-3) and brand compliance verification (Task 5-1-4). Lower scores in both Task 5-1-3 and Task 5-1-4 highlight a broader weakness in the ability to accurately compare key information, a core requirement for both tasks. This difficulty is then compounded by the unique demands of each: the former requires complex reasoning and calculation, while the latter necessitates the accurate parsing of tables in Markdown format. In contrast, performance on the information extraction task (Task 5-1-5) showcases the LLMs’ proficiency in parsing and extracting information from long-context domain documents.
In the Evaluation tasks, the modest Kendall’s tau coefficients (failing to surpass the 0.7 threshold for strong correlation) are attributable to the profound difficulty of shifting from simple pattern matching to the kind of high-fidelity, evidence-based reasoning process. This process requires LLMs to interpret key evidence in criteria that may be paraphrased, expressed implicitly, or scattered across a long context, posing a significant challenge to the goal of substituting human experts in rubric-based evaluations. In the Creation tasks, all models failed to surpass the 60-point threshold. The uniformly limited performance is attributable to the complex demands of professional document generation. This process requires a multi-faceted synthesis of factual accuracy, logical coherence, and structural norms, posing a significant challenge to the models’ advanced generative capabilities.
5 Discussion
5.1 Systematic bias of LLM-as-a-Judge approach
Figure 10(a) presents a scatter plot of LLM prediction scores against human expert scores, where the size of each point is proportional to its frequency of occurrence. A y=x line in this figure indicates the ground truth for perfect model-human agreement. For each model, a smoothed curve is fitted to the data points to illustrate its performance trend. While all models demonstrate a generally positive correlation with human assessments, as shown by their monotonically rising curves, DeepSeek-R1’s curve is closest to the y=x line, indicating the highest level of agreement.
However, the plot also reveals a systematic bias: for instance, models tend to assign an average score of 4.25 to documents rated in the 0–2 range by experts, while conversely assigning an average of only 7.96 to those rated between 8 and 10. This tendency to overestimate low scores and underestimate high scores results in a compressed scoring interval. This compression has significant implications for using any model in automated grading, particularly those with less-than-ideal performance. Therefore, to meet the required standards of reliability, it is essential to introduce calibration methods and uncertainty assessments.
Two calibration methods (i.e., isotonic regression and piecewise linear regression) were applied in this paper to mitigate the observed scoring bias [50, 51]. As illustrated in Figure 10(b), both methods successfully shifted the model’s performance curve closer to the ideal y=x line. This improvement is quantitatively confirmed by a significant reduction in the mean absolute error (MAE) from an initial 2.947 to 1.926 (isotonic regression) and 2.015 (piecewise linear regression). These results demonstrate that both methods effectively reduce the systematic bias of LLM-as-a-Judge, leading to predictions that align more closely with human expert scores.
Figure 10: Scatter Plots with LOWESS Curves of Model Performance in Evaluation Tasks
(a) original performance of LLMs; (b) performance of DeepSeek-V3 with calibration methods
5.2 Performance decline in tasks related to code tables
Figure 11 presents model performance in two types of comparative tasks designed to specifically assess the model’s ability to understand and reason over knowledge in code tables. A significant and consistent performance degradation on these two tasks could be clearly observed. To investigate the reasons for this decline in consistency, this paper provides the model with the necessary knowledge sources in two distinct approaches. These two approaches enable LLMs to generate responses based on tabular content embedded in codes, as delineated in Figure 12. Approach #1 converts tabular content into natural language context through expert-crafted textual descriptions. Approach #2 utilizes models with multi-modal capabilities (GPT-4.1 in this paper) to automatically transform image-based tables into HTML formats, which is less intuitive for human interpretation but inherently machine-parsable [52]. Table 3 shows the performance of models on these two tasks. Both Approach #1 and Approach #2 achieved significant improvements over the baseline results (’Original’). The average accuracies on these two tasks improved from a baseline of 45.27% and 40.65% to 98.94% and 87.27% with Approach #1, and to 72.67% and 68.24% with Approach #2, respectively. These significant improvements also suggest that the decline stems from the model’s inability to properly encode knowledge from the code tables.
Figure 11: Performance decline in questions related to tables in codes
Figure 12: Two context formatting approaches for tables in codes
Moreover, this paper compares two effective approaches for transforming knowledge within tables. After being processed by these approaches, the tables can not only serve as context for the model, as shown in Figure 12, but also have their knowledge converted into a pre-training dataset for the model. The superior performance of Approach #1 over Approach #2 can be attributed to Approach #2’s inability to correctly reproduce the complex layout of tables from building codes, particularly those with spanning cells. Even with advanced multi-modal architectures or dedicated parsing methods, lossless conversion of tabular content remains unattainable [53, 54, 55]. This inherent limitation underscores a fundamental trade-off in transforming tabular data for LLMs: while manual, expert-driven methods can achieve high fidelity, they lack scalability; conversely, automated approaches, though rapid and scalable, often struggle to preserve the complete accuracy and integrity of the original content.
Table 3: Performance comparison of two context formatting approaches
Model Task 2-3 Task 3-3
Original Appr. 1 Appr. 2 Original Appr. 1 Appr. 2
Moonshot-v1-128k 43.92 97.88(+53.96) 71.96(+28.04) 34.65 80.71(+46.06) 62.99(+28.34)
GLM-4-Plus 34.92 100(+65.08) 70.9(+35.98) 35.83 87.01(+51.18) 66.54(+30.71)
Qwen-Turbo 43.92 98.41(+54.49) 68.78(+24.86) 43.31 83.07(+39.76) 64.57(+21.26)
QwQ-32B 40.74 98.41(+57.67) 73.54(+32.8) 42.13 88.19(+46.06) 71.65(+29.52)
DeepSeek-V3 52.38 100(+47.62) 68.78(+16.4) 32.68 84.25(+51.57) 60.24(+27.56)
DeepSeek-R1 63.49 99.47(+35.98) 76.72(+13.23) 53.94 90.94(+37) 74.41(+20.47)
GPT-4o 44.44 97.88(+53.44) 74.6(+30.16) 39.76 91.73(+51.97) 72.83(+33.07)
GPT o3-mini 39.68 100(+60.32) 75.66(+35.98) 42.91 90.55(+47.64) 70.87(+27.96)
Hunyuan-TurboS 43.92 98.41(+54.49) 73.02(+29.1) 44.49 88.98(+44.49) 70.08(+25.59)
6 Conclusion
To guide the responsible integration and future development of large language models (LLMs) in augmenting the complex, knowledge-intensive tasks of the Architecture, Engineering, and Construction (AEC) industry, this paper established a foundational evaluation benchmark, AECBench, to assess the strengths and limitations of current LLMs in these contexts. A five-level hierarchical framework (Knowledge Memorization, Knowledge Understanding, Knowledge Reasoning, Knowledge Calculation, and Knowledge Application) was established to mirror domain-specific cognitive demands. A high-quality dataset of 4,800 questions across 23 tasks was primarily created by engineers and subsequently validated by domain experts through a two-round review process. A scalable, LLM-driven evaluation pipeline guided by expert-defined rubrics was developed to assess the complex and long-form responses to open-ended questions.
The evaluation of nine LLMs revealed a clear performance trend of LLMs in AEC applications. While the models demonstrate exceptional proficiency in memorizing domain-specific knowledge, their performance consistently degrades as the cognitive complexity of the tasks increases. This trend reveals several fundamental limitations of current LLMs in the AEC domain: (i) a notable deficiency in complex reasoning and calculation; (ii) a modest correlation with human expert judgments in rubric-based evaluations, and (iii) a lack of factual accuracy in professional document generation. DeepSeek-R1 demonstrates strong performance, consistently placing it among the top-performing models in AECBench.
The models’ capabilities as rubric-based judges were thoroughly discussed in this paper. The analysis uncovers a systematic bias, and in response, two effective calibration methods (i.e., isotonic regression and piecewise linear regression) were proposed to enhance their reliability. The study also discussed the models’ unexpectedly low performance on tasks involving tabular content from building codes. The low performance reveals another limitation about an insufficient mastery of the knowledge within building code tables. While two viable context formatting approaches (i.e., manual textual description and automated HTML conversion) were established, the challenge of efficiently transforming tabular data into fully utilizable and correct content for LLMs remains a critical area for future work.
Future work would incorporate tasks related to multimodal data, such as architectural drawing recognition, to better reflect the full spectrum of information modalities in the AEC field. It is noted that AECBench will be made open-source shortly afterwards. The community is invited to contribute to the enhancement.
Acknowledgment
This project is supported by Shanghai Qi Zhi Institute (Grant No.SQZ202309) and the Scientific Research Project Fund of Arcplus Group PLC (24-1L-0142-Z). We extend our gratitude to every engineer and researcher who contributed to this dataset.
Appendix A Details of tasks
A.1 Knowledge Memorization
Figure 13: The instruction and an example of Task 1-1 Code Memorization & Retrieval
Figure 14: The instruction and an example of Task 1-2 Terminology Memorization
Figure 15: The instruction and an example of Task 1-3 Abbreviation Memorization
A.2 Knowledge Understanding
Figure 16: The instruction and an example of Task 2-1 Code Provision Interpretation
Figure 17: The instruction and an example of Task 2-2 Design Literacy Q & A
Figure 18: The instruction and an example of Task 2-3 Code Provision Interpretation(Tabular Data)
A.3 Knowledge Reasoning
Figure 19: The instruction and an example of Task 3-1 Design Decision Formulation
Figure 20: The instruction and an example of Task 3-2 Building Type Prediction
Figure 21: The instruction and an example of Task 3-3 Design Decision Formulation(Tabular Data)
A.4 Knowledge Calculation
Figure 22: The instruction and an example of Task 4-1 Calculation in Architectural Design
Figure 23: The instruction and an example of Task 4-2 Calculation in MEP Design
Figure 24: The instruction and an example of Task 4-3 Calculation in Construction
Figure 25: The instruction and an example of Task 4-4 Calculation in Engineering Economics
A.5 Knowledge Application
Figure 26: The instruction and an example of Task 5-1-1 Document Classification
Figure 27: The instruction and an example of Task 5-1-2 Document Proofreading
Figure 28: The instruction and an example of Task 5-1-3 Compliance Checking
Figure 29: The instruction and an example of Task 5-1-4 Brand Compliance Verification
Figure 30: The instruction and an example of Task 5-1-5 Information Extraction (continued on next page)
Figure 30: The instruction and an example of Task 5-1-5 Information Extraction (continued)
Figure 30 : The instruction and an example of Task 5-1-5 Information Extraction (continued)
Figure 31: The instruction and an example of Task 5-2-1 Evaluation of Documents in Architectural Design
(continued on next page)
Figure 31: The instruction and an example of Task 5-2-1 Evaluation of Documents in Architectural Design(continued)
Figure 31: The instruction and an example of Task 5-2-1 Evaluation of Documents in Architectural Design(continued)
Figure 32: The instruction and an example of Task 5-2-2 Evaluation of Documents in Structural Design
(continued on next page)
Figure 32: The instruction and an example of Task 5-2-2 Evaluation of Documents in Structural Design(continued)
Figure 32: The instruction and an example of Task 5-2-2 Evaluation of Documents in Structural Design(continued)
Figure 33: The instruction and an example of Task 5-3-1 Conceptual Design Proposal Generation
Figure 34: The instruction and an example of Task 5-3-2 Specialized Report Generation
Figure 35: The instruction and an example of Task 5-3-3 Tender Review Report Generation